OpenStudy (anonymous):

Choose the equation for the line passing through (4,3) and parallel to 4x-5y=10?

OpenStudy (anonymous):

line parllel to 4x-5y=10 is 4x-5y=k (where k is constant to be determined) since the line passes through4,3 thus putting for x=4 and y=3 in 4x-5y=k we have k=16-15=1 thus the required line is 4x-5y=1

OpenStudy (anonymous):

cool thanks that really helped